# Locus: Agentic Predicate Synthesis for Directed Fuzzing

### v1

@@ 1. Introduction
DGF is challenging, as the target program states are often deeply nested in the program, while the search space introduced by the complexity of real-world software is prohibitively large. To speed up the search and schedule promising inputs, most existing techniques leverage the control flow proximity, e.g., the distance to target location in the control flow graph, or heuristics based on the semantics of the branch predicates (Chen and Chen, 2018; Li et al., 2017; Shah et al., 2022), e.g., the target state if x==42 has the distance metric of |x42||x-42|. However, such feedback is sometimes too sparse or indirect to reliably measure the progress, especially when there is a long chain of implicit preconditions guarding the target program states (Mitropoulos et al., 2023; Holler et al., 2012; Dutra et al., 2023; Rigger and Su, 2020; Klees et al., 2018; Zhang et al., 2023; Kim et al., 2023; Sherman and Nagy, 2025). For example, triggering CVE-2018-13785 in libpng requires a PNG file to satisfy a precise sequence of preconditions, i.e., valid signature, correct chunk layout, specific IHDR fields (e.g., bit depth, color type), and a magic image width (0x55555555), to trigger an integer overflow (Kim et al., 2023), while the predicates to explicitly check these conditions are largely absent in the code to provide an incremental progress guidance.
@@
To capture the intricate feedback to improve search efficiency, more advanced approaches identify progress-capturing constraints in the program to drive execution towards satisfying specific temporal orders and preconditions (Meng et al., 2022; Holler et al., 2012; Lee et al., 2021; Liang et al., 2020; Li et al., 2017; Fioraldi et al., 2021; Padhye et al., 2019; Ispoglou et al., 2020). However, these constraints are often manually crafted by experts and tailored to a few specific target state types, e.g., focusing on a temporal memory safety bug like a use-after-free by enforcing an allocate–free–use sequence. As different programs can have diverse target states and disparate functionalities, the feedback metric effective in one case may not generalize to another (Chen et al., 2019).
@@
As Machine Learning (ML) and Large Language Models (LLMs) have demonstrated surprising code reasoning capabilities, there has been a growing interest in extending such capabilities to help guide fuzzing (Patra and Pradel, 2016). A common approach is to employ LLMs to directly generate inputs or grammar-aware input generators (i.e., fuzz driver or harness) (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d; Lyu et al., 2024; Chen et al., 2025; Yang et al., 2025a), where the preconditions useful to reach target states are expressed as part of the input grammar constraints. However, not all conditions to reach target states can be easily represented as input grammars. For example, Figure 1 shows that an intermediate state (!found_plte) necessary to trigger the buffer overflow in libpng only emerges in the middle of the execution and cannot be easily checked at the input level. More importantly, such a task formulation is particularly challenging for LLMs, as reasoning from the target program states all the way to the input often requires an extremely long context and convoluted reasoning chain (Jiang et al., 2024; Ding et al., 2024a; Li et al., 2024a; Risse et al., 2025; Risse and Böhme, 2024), a common pitfall for LLM hallucination (Sahoo et al., 2024; Yao et al., 2025), let alone the challenge to verify the correctness of the LLM generated inputs and harnesses, which can significantly impede the fuzzing progress if the generated constraints are incorrect (Padhye et al., 2019; Ispoglou et al., 2020).
@@ 5. Related Work
Direct grey-box fuzzing  Directed grey-box fuzzing is particularly challenging, primarily due to the prohibitively large search space with sparse rewards(Li et al., 2024b; Luo et al., 2023), i.e., unlike undirected fuzzing where any new coverage indicates progress. Most prior works develop heuristics through static analysis by computing path distances to target states to narrow down the search space (She et al., 2024; Shah et al., 2022; Huang et al., 2024a; Böhme et al., 2017; Luo et al., 2023; Srivastava et al., 2022b; Li et al., 2024b; Chen and Chen, 2018; Lee et al., 2021). For example, AFLGo(Böhme et al., 2017) introduces the idea of using distance metrics between test inputs and target basic blocks to prioritize seeds that are closer to the target, while CAFL(Lee et al., 2021) further improves this metric by introducing specialized progress-capturing state representation and calculating the distance from the testing inputs to the nearest state.

### v2

@@ 1. Introduction
DGF is challenging, as the target program states are often deeply nested in the program, while the search space introduced by the complexity of real-world software is prohibitively large. To speed up the search and schedule promising inputs, most existing techniques leverage the control flow proximity, e.g., the distance to target location in the control flow graph, or heuristics based on the semantics of the branch predicates (Chen and Chen, 2018; Li et al., 2017; Shah et al., 2022), e.g., the target state if x==42 has the distance metric of |x42||x-42|. However, such feedback is sometimes too sparse or indirect to reliably measure the progress, especially when there is a long chain of implicit preconditions guarding the target program states (Mitropoulos et al., 2023; Holler et al., 2012; Dutra et al., 2023; Rigger and Su, 2020; Klees et al., 2018; Zhang et al., 2023; Kim et al., 2023; Sherman and Nagy, 2025; Aschermann et al., 2020). For example, triggering CVE-2018-13785 in libpng requires a PNG file to satisfy a precise sequence of preconditions, i.e., valid signature, correct chunk layout, specific IHDR fields (e.g., bit depth, color type), and a magic image width (0x55555555), to trigger an integer overflow (Kim et al., 2023), while the predicates to explicitly check these conditions are largely absent in the code to provide an incremental progress guidance.
@@
To capture the intricate feedback to improve search efficiency, more advanced approaches identify progress-capturing constraints in the program to drive execution towards satisfying specific temporal orders and preconditions (Meng et al., 2022; Holler et al., 2012; Lee et al., 2021; Liang et al., 2020; Li et al., 2017; Fioraldi et al., 2021; Padhye et al., 2019; Ispoglou et al., 2020; Balzarotti, 2021; Aschermann et al., 2020). However, these constraints are often manually crafted by experts and tailored to a few specific target state types, e.g., focusing on a temporal memory safety bug like a use-after-free by enforcing an allocate–free–use sequence. As different programs can have diverse target states and disparate functionalities, the feedback metric effective in one case may not generalize to another (Chen et al., 2019).
@@
As Machine Learning (ML) and Large Language Models (LLMs) have demonstrated surprising code reasoning capabilities, there has been a growing interest in extending such capabilities to help guide fuzzing (Patra and Pradel, 2016). A common approach is to employ LLMs to directly generate inputs or grammar-aware input generators (i.e., fuzz driver or harness) (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d; Lyu et al., 2024; Chen et al., 2025; Yang et al., 2025a), where the preconditions useful to reach target states are expressed as part of the input grammar constraints. However, not all conditions to reach target states can be easily represented as input grammars. For example, Figure 1 shows that an intermediate state (!found_plte) necessary to trigger the buffer overflow in libpng only emerges in the middle of the execution and cannot be easily checked at the input level. More importantly, such a task formulation is particularly challenging for LLMs, as reasoning from the target program states all the way to the input often requires an extremely long context and convoluted reasoning chain (Jiang et al., 2024; Li et al., 2025; Ding et al., 2024a; Li et al., 2024a; Risse et al., 2025; Risse and Böhme, 2024; Li et al., 2024b; Risse et al., 2025), a common pitfall for LLM hallucination (Sahoo et al., 2024; Yao et al., 2025), let alone the challenge to verify the correctness of the LLM generated inputs and harnesses, which can significantly impede the fuzzing progress if the generated constraints are incorrect (Padhye et al., 2019; Ispoglou et al., 2020).
@@ 5. Related Work
Direct grey-box fuzzing  Directed grey-box fuzzing is particularly challenging, primarily due to the prohibitively large search space with sparse rewards(Li et al., 2024c; Luo et al., 2023), i.e., unlike undirected fuzzing where any new coverage indicates progress. Most prior works develop heuristics through static analysis by computing path distances to target states to narrow down the search space (She et al., 2024; Shah et al., 2022; Huang et al., 2024a; Böhme et al., 2017; Luo et al., 2023; Srivastava et al., 2022b; Li et al., 2024c; Chen and Chen, 2018; Lee et al., 2021). For example, AFLGo(Böhme et al., 2017) introduces the idea of using distance metrics between test inputs and target basic blocks to prioritize seeds that are closer to the target, while CAFL(Lee et al., 2021) further improves this metric by introducing specialized progress-capturing state representation and calculating the distance from the testing inputs to the nearest state.